OpenStudy (babybritt002):

0.005kg = ___dag Please use factor label method

OpenStudy (anonymous):

1kg=100 decagram dag=decagram \[0.005kg \times \frac{ 100 decagram}{ 1kg }=\]

OpenStudy (babybritt002):

how is 1kg=100 decagram I thought i would be 1kg=10decagram

OpenStudy (mathstudent55):

1 kg = 1000 g 1 dag = 10 g \(0.005 ~kg \times \dfrac{1000~g}{1~kg} \times \dfrac{1~dag}{10~g}\) \(=0.005 ~\cancel{kg} \times \dfrac{1000~\cancel{g}}{1~\cancel{kg}} \times \dfrac{1~dag}{10~\cancel{g}}\) \(= \dfrac{ 0.005 \times 1000~ }{10~}~dag\)

OpenStudy (mathstudent55):

A kilogram is 1000 g A decagram is 10 g That means 1 kg is 100 dag

OpenStudy (anonymous):

Well its not.

OpenStudy (mathstudent55):

Since you normally don't relate kg to dag, but you do relate kg to g and dag to g, I think it is better to do these conversions going through grams.

OpenStudy (anonymous):

\[da=10^1\] \[kg=10^3\] \[1kg=\frac{ 10^3 }{ 10^1 }=10^2=100\]

OpenStudy (mathstudent55):

1 hg (hectogram) = 100 g, so 1 kg = 10 hg. 1 dag (decagram) = 10 g, so 1 kg = 100 dag.

hero (hero):